OpenStudy (anonymous):

if i am adding fractions and the first one is a negative do i switch the fractions around and put the negative fraction at the back so i can just subtract

OpenStudy (anonymous):

i would

OpenStudy (chaise):

You can do that, if it makes it easier for you :)

OpenStudy (anonymous):

but would i lose marks for doing that?

OpenStudy (anonymous):

some math teacher love starting with negatives \[-\frac{2}{5}+\frac{1}{2}\]

OpenStudy (anonymous):

but are we allowed to switch them in exams?

OpenStudy (anonymous):

i would write \[\frac{1}{2}-\frac{2}{5}\]

OpenStudy (anonymous):

but how do we solve wothout switching

OpenStudy (anonymous):

i cannot answer that, because i don't know what your teacher expects. but the truth is you are going to end up subtracting anyway

OpenStudy (anonymous):

can u show me how to do it without switching

OpenStudy (chaise):

Basicly you have to imagine a number lime. -(2/5)+(1/2) -0.4+0.5 If we are 0.4 spaces away from 0, and we add 0.5, we get 0.1 spaces infront of zero. 0.1 can be written as 1/10.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

ooooooooooooooooooooooooooooooooooooooooooooooooooooh kay iGOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOT YEPPPPPPPPPPPPPPPPPPIEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE